OpenStudy (anonymous):

3^{x+3}-9^{2x-1}=0 what are the steps to finding the solution?

OpenStudy (kamille):

firstly, you need to change 9 to 3^2

OpenStudy (kamille):

What will you get?

OpenStudy (anonymous):

3^{x+3} = 3^2^{2x-1}=0

OpenStudy (anonymous):

ignore = 0

OpenStudy (campbell_st):

rewrite the equation in terms of base 3 and then equate the 2 terms \[3^{x + 3}= 3^{2(2x - 1)}\] now they are the same, just solve x + 3 = 2(2x -1) for x.

OpenStudy (kamille):

So, now you can say that x+3=2(2x+1) and you need to find x now

OpenStudy (anonymous):

how do i solve for x

OpenStudy (kamille):

x+3=4x+2 now you need to movie all numbers which have x to left, and all numbers which don't have x to the right. try it.

OpenStudy (anonymous):

cause its -1 not +1 it looks like x+3=4x+2 then 5x=-1??

OpenStudy (kamille):

Oh,sorry, my bad. x+3=2(2x-1) x+3=4x-2 -3x=-5|:(-3) x=1 2/3

OpenStudy (anonymous):

5|:(-3)???

OpenStudy (anonymous):

how does x = 12/3?

OpenStudy (anonymous):

x should equal 1.667 cause (-5)/(-3)

myininaya (myininaya):

He wrote it as a mixed fraction. Also in the future, please don't post the same question twice. You should either bump or get more eyes by posting a link to your question in the math chat. Thanks.